# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

### v1

@@ 4 EquiReg: Equivariance Regularized Diffusion
Table 1: Effect of EquiReg on ReSample [37]. Inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.01$.
@@
Super-res. (×$\times 4$) ReSample 0.204 40.46 28.02 0.790
@@
We show that models trained on PDEs with inherent symmetries can indeed learn the equivariance property. Neural operators [130] are a popular architecture for modeling physics systems, as their function space setup can preserve the continuum of physics functions and generalize well to different resolutions. Among these, the Fourier Neural Operator (FNO [131]) has drawn the most attention. We train an FNO on the inverse problem of Navier-Stokes, which respects reflection symmetry (see details in Appendix), without explicit data augmentation. As shown in Figure 4(b), FNOs trained on non-augmented physics data yield lower equivariance error ($\text{Equi}_{\text{pixel}}$ loss: $\mathcal{R}({\bm{x}}_{t})=\|S_{g}(\text{FNO}({\bm{x}}_{0|t}))-\text{FNO}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$) on in-distribution as opposed to out-of-distribution data.
5 Results
This section provides experimental results on the performance of EquiReg for inverse problems, including linear and nonlinear image restoration tasks and reconstructing PDEs. To explore the impact of scaling the noise scheduler, the Appendix includes an experiment on DreamSampler [82], showing preliminary evidence that EquiReg can improve performance on text-to-image guidance.
@@ 5 Results
Experimental setup for image restorations. We evaluate the performance of EquiReg when applied to the following three diffusion-based inverse solvers: PSLD [38], ReSample [37], and DPS [32]. Of these methods, PSLD and ReSample are latent solvers, while DPS is a pixel-based solver. We compare against several other methods including DDRM [34], MCG [33], DiffPIR [39], MPGD-AE [62], and DiffStateGrad [41]. We measure performance via perceptual similarity (LPIPS), distribution alignment (FID), pixel-wise fidelity (PSNR), and structural consistency (SSIM). We test EquiReg on two datasets: a) the FFHQ $256\times 256$ validation dataset [132], and b) the ImageNet $256\times 256$ validation dataset [133]. For pixel-based experiments, we use i) the pre-trained diffusion model from [32] on the FFHQ dataset, and ii) the pre-trained model from [68] on the ImageNet dataset. For latent diffusion experiments, we use i) the unconditional LDM-VQ-4 model [61] on FFHQ, and ii) the Stable Diffusion v1.5 [61] model on ImageNet.
Table 2: Linear inverse problems on FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
Method Gaussian deblur Motion deblur Super-resolution (×$\times 4$)
[…]
Latent
PSLD 0.357 106.2 22.87 0.537 0.322 84.62 24.25 0.615 0.313 89.72 24.51 0.650
Equi-PSLD (ours) 0.344 94.09 24.42 0.657 0.338 99.14 24.83 0.669 0.289 90.88 26.32 0.718
EquiCon-PSLD (ours) 0.320 83.18 24.38 0.664 0.322 89.87 25.14 0.699 0.277 79.39 26.14 0.722
Pixel-based
DDRM 0.239 92.43 24.93 0.732 - - - - 0.282 79.25 26.58 0.782
MCG 0.340 101.2 6.72 0.051 0.702 310.5 6.72 0.055 0.520 87.64 20.05 0.559
DiffPIR 0.236 59.65 27.36 - 0.255 65.78 26.57 - 0.260 65.77 26.64 -
MPGD-AE 0.150 114.9 24.42 0.664 0.120 104.5 25.72 0.723 0.168 137.7 24.01 0.670
DPS 0.145 104.8 25.48 0.709 0.132 99.75 26.75 0.754 0.191 125.4 24.38 0.675
DiffStateGrad-DPS 0.128 52.73 26.29 0.739 0.118 50.14 27.61 0.782 0.186 73.02 24.65 0.683
Equi-DPS (ours) 0.114 48.76 26.32 0.750 0.094 41.71 28.23 0.810 0.120 51.00 27.15 0.776
(a) Deblurring and super-resolution linear inverse problems.
Method Box inpainting Random inpainting
[…]
Latent
PSLD 0.158 43.02 24.22 0.819 0.246 49.77 29.05 0.809
Equi-PSLD (ours) 0.098 31.54 24.19 0.875 0.188 41.61 30.43 0.878
EquiCon-PSLD (ours) 0.092 35.07 24.26 0.884 0.204 40.75 29.99 0.847
Pixel-based
DDRM 0.207 78.62 22.26 0.801 0.218 - 25.75 -
MCG 0.309 40.11 19.97 0.703 0.286 29.26 21.57 0.751
MPGD-AE 0.138 248.7 21.59 0.753 0.172 339.0 25.22 0.732
DPS 0.133 56.89 23.10 0.799 0.113 51.32 29.63 0.849
DiffStateGrad-DPS 0.114 47.53 24.10 0.831 0.107 49.42 30.15 0.852
Equi-DPS (ours) 0.099 40.47 23.39 0.835 0.068 33.65 32.16 0.902
(b) Inpainting linear inverse problems.
Figure 5: EquiReg consistently improves performance across a range of measurement noise levels. Gaussian deblur using FFHQ $256\times 256$ for $\sigma_{{\bm{y}}}=[0,0.8]$.
We evaluate EquiReg on a variety of linear and nonlinear restoration tasks for natural images. We fix sets of $100$ images from FFHQ and ImageNet as our validation sets. All images are normalized from $[0,1]$. For the majority of experiments, we use noise level $\sigma_{{\bm{y}}}=0.05$ (we indicate $\sigma_{{\bm{y}}}$ in our tables). For linear inverse problems, we consider (1) box inpainting, (2) random inpainting, (3) Gaussian deblur, (4) motion deblur, and (5) super-resolution. We apply a random $128\times 128$ pixel box for box inpainting, and a $70\%$ random mask for random inpainting. For Gaussian and motion deblur, we use kernels of size $61\times 61$, with standard deviations of $3.0$ and $0.5$, respectively. For super-resolution, we downscale images by a factor of $4$ using a bicubic resizer. For nonlinear inverse problems, we consider (1) phase retrieval, (2) nonlinear deblur, and (3) high dynamic range (HDR). We use an oversampling rate of $2.0$ for phase retrieval, and due to instability of the task, we generate four independent reconstructions and take the best result (as also done in DPS [32], DAPS [40], and DiffStateGrad [41]). We use the default setting from [134] for nonlinear deblur, and a scale factor of 2 for HDR.
Table 3: Linear inverse problems on ImageNet 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.05$.
[…]
PSLD 263.9 20.70 252.1 21.26 224.3 22.29 151.4 16.28 83.22 26.56
EquiCon-PSLD (ours) 214.5 22.01 196.3 22.69 198.5 22.34 137.6 19.25 65.14 27.03
Figure 6: EquiReg reduces failure cases and enhances reconstruction fidelity. Super-resolution using FFHQ $256\times 256$.
@@
For these tasks, we adopt the pre-trained encoder-decoder $\mathcal{E}$-$\mathcal{D}$ as our MPE function. For FFHQ, the symmetry group contains reflection (across vertical axis). For ImageNet, we define a rotation group $G=\{\nicefrac{{\pi}}{{2}},\nicefrac{{3\pi}}{{2}}\}$, and uniformly at random select the group action for each sample. Finally, the loss functions given in Equation 6 are used as regularization terms.
Image restoration performance. Our results (Tables 1, 2 and 3) show that EquiReg improves the performance of several methods across a wide variety of linear and nonlinear tasks on both FFHQ and ImageNet. EquiReg significantly improves performance on box inpainting when applied to ReSample (Table 1). We attribute this improvement in part to the reduction of failure cases (Figure 6).
@@
We observe that EquiReg achieves its largest improvements on perceptual metrics (FID and LPIPS), suggesting that it generates more realistic images that lie closer to the data manifold (see Appendix for supporting qualitative results). Moreover, EquiReg improves performance under high measurement noise (Figure 5). This result aligns with Figure 4(a), which shows the equivariance error is lower on clean images than noisy ones, indicating that EquiReg enforces an effective denoising.
PDE reconstruction tasks. We evaluate the performance of EquiReg on two important PDE problems: the Helmholtz and Navier-Stokes equations (see Appendix). The objective is to solve both forward and inverse problems in sparse sensor settings. The forward problem involves predicting the solution function or the final state using measurements from only $3\%$ of the coefficient field or the initial state. The inverse problem, conversely, aims to predict the input conditions from observations of $3\%$ of the system’s output. This task is challenging due to the nonlinearity of the equations, the complex structure of Gaussian random fields, and the sparsity of observations.
Table 4: Partially-observed PDE reconstruction.
@@
Recent studies [51, 56, 57] have demonstrated the superiority of diffusion-based approaches over deterministic single-forward methods for PDE reconstructions. DiffusionPDE [51] decomposes the conditional log-likelihood into a data prior (from a pre-trained diffusion model) and a measurement score, thereby providing solutions via conditional sampling. FunDPS [56, 57] further extends the sampling process to a more natural infinite-dimensional spaces, achieving better accuracy and speed via function space models.
@@
We integrate EquiReg into the state-of-the-art FunDPS framework [56, 57], computing the Equi loss using data inherent equivariance derived from an FNO trained on the corresponding inverse problem. For our experiments, we use reflection symmetry (i.e., flipping along the $y=x$ axis), and observe no significant performance difference when using other transformations such as rotations or alternating flips. As shown in Table 4, Equi-FunDPS improves performance, measured by relative $\ell_{2}$ loss—across various tasks, especially in inverse problems where a strong data prior is critical.
@@ 6 Conclusion
We introduce Equivariance Regularized (EquiReg) diffusion for solving inverse problems. EquiReg corrects errors introduced in posterior approximations by regularizing the sampling trajectories so they lie closer to the data manifold. Our method applies to many inverse problem settings across diverse domains and is flexible to the choice of diffusion model and sampling method.
Limitations and future work. EquiReg’s effectiveness relies on the quality of the pre-trained diffusion prior. In addition, we do not address the approximations of the underlying diffusion models, but we regularize them for improved performance. Moreover, applying EquiReg requires task-specific regularizer design: one must choose an appropriate symmetry group and identify suitable MPE functions for each problem setting. As a result of this task-specific design, EquiReg can be explored in many other problem settings where equivariance is training induced or data inherent. We leave this for future work.
7 Acknowledgements and Contributions
Aditi worked on this project as part of her bachelor’s thesis. Anima was supported by Bren Chair Professor and AI2050 Senior Fellow.
Bahareh proposed and led the project. Aditi and Rayhan implemented the method and executed all the image restoration experiments. Abbas conducted the text-to-image experiments, and Jiachen performed the PDE reconstruction experiments. Bahareh wrote the original draft of the paper with Aditi, Rayhan, and Jiachen writing the results section. The project was originally based on training-induced MPE functions, proposed by Bahareh. Later versions of the project include data-inherent MPE functions, proposed by Jiachen; this is used for the PDE experiments. Bahareh, Abbas, Jiachen, and Chuwei collaborated on formalizing the theory, and the final formulation of Wasserstein gradient flow was proposed and proven by Chuwei. All authors provided editorial comments and feedback on the manuscript. Anima supervised the project and provided valuable feedback.
Appendices
@@ 7 Acknowledgements and Contributions
We provide our source code when EquiReg is implemented on DPS. This supplementary materials contain the following:
@@
  • Appendix A includes additional experiments on text-to-image guidance. We regularize DreamSampler [82] with EquiReg for an improved performance (see Figures 7, 8, 9 and 10).
  • Appendix B includes additional image restoration experiments for ImageNet. We show that EquiReg improves LatentDAPS and DPS across several different inverse problems (Table 5, Table 6).
  • Appendix C reports standard deviations on experiments done on PSLD (see Table 7).
@@
  • Appendix D includes qualitative analysis on the performance of methods with and without EquiReg. Results show a reduction of artifacts and an improved perceptual quality of the solution. This section also includes the equivariance error of a pre-trained encoder used in EquiReg (Figure 11(a)), an equivariant loss vs. Gaussian noise plot (Figure 12), and a histogram of Equi’s improvement on DPS (Figure 13).
@@
  • Appendix E demonstrates EquiReg implementation for PSLD, ReSample, DPS, and LatentDAPS (Algorithm 2, Algorithm 3, Algorithm 4, Algorithm 5, Algorithm 6, Algorithm 7). It also contains information about the EquiReg hyperparameters for image restoration tasks.
@@
  • Appendix F contains information on the PDE reconstruction experiment. It discusses the equations along with implementation details and hyperparameters.
  • Appendix G provides theoretical proof of Propositions 4.1 and 4.2.
@@
  • Appendix H contains additional background information on solving inverse problems, vanishing-error autoencoders, and equivariance.
  • Appendix I discloses computing resources used to conduct the experiments.
  • Appendix J credits code assets used for our experiments.
  • Appendix K discusses the broader impacts of this paper, the developed method, the conducted experiments, and their overall implications.
  • Appendix L concludes the appendix with a “responsible release” statement.

### v2

@@ 4 Results
Table 3: EquiReg for diffusion models on FFHQ. $256\times 256$ with $\sigma_{{\bm{y}}}=0.05$.
Method Gaussian deblur Motion deblur Super-resolution ($\times 4$) Box inpainting Random inpainting
[…]
PSLD 0.357 106.2 22.87 0.322 84.62 24.25 0.313 89.72 24.51 0.158 43.02 24.22 0.246 49.77 29.05
Equi-PSLD 0.344 94.09 24.42 0.338 99.14 24.83 0.289 90.88 26.32 0.098 31.54 24.19 0.188 41.61 30.43
EquiCon-PSLD 0.320 83.18 24.38 0.322 89.87 25.14 0.277 79.39 26.14 0.092 35.07 24.26 0.204 40.75 29.99
(a) Latent diffusion.
Method Gaussian deblur Motion deblur Super-resolution ($\times 4$) Box inpainting Random inpainting
[…]
DPS 0.145 104.8 25.48 0.132 99.75 26.75 0.191 125.4 24.38 0.133 56.89 23.10 0.113 51.32 29.63
Equi-DPS (ours) 0.114 48.76 26.32 0.094 41.71 28.23 0.120 51.00 27.15 0.099 40.47 23.39 0.068 33.65 32.16
DiffStateGrad-DPS 0.128 52.73 26.29 0.118 50.14 27.61 0.186 73.02 24.65 0.114 47.53 24.10 0.107 49.42 30.15
MCG 0.340 101.2 6.72 0.702 310.5 6.72 0.520 87.64 20.05 0.309 40.11 19.97 0.286 29.26 21.57
MPGD-AE 0.150 114.9 24.42 0.120 104.5 25.72 0.168 137.7 24.01 0.138 248.7 21.59 0.172 339.0 25.22
(b) Pixel-based diffusion.
Table 4: EquiReg for latent diffusion models on ImageNet. $256\times 256$ with $\sigma_{{\bm{y}}}=0.05$.
[…]
PSLD 263.9 20.70 252.1 21.26 224.3 22.29 151.4 16.28 83.22 26.56
EquiCon-PSLD 214.5 22.01 196.3 22.69 198.5 22.34 137.6 19.25 65.14 27.03
Table 5: EquiReg for ReSample on linear and nonlinear tasks. FFHQ 256 $\times$ 256 with $\sigma_{{\bm{y}}}=0.01$.
@@
Super-res. ($\times 4$) ReSample 0.204 40.46 28.02 0.790
@@
We measure performance via perceptual similarity (LPIPS), distribution alignment (FID), pixel-wise fidelity (PSNR), and structural consistency (SSIM). We test EquiReg on a) the FFHQ $256\times 256$ (Karras et al., 2021) and b) ImageNet $256\times 256$ validation set (Deng et al., 2009). For pixel-based experiments, we use the pre-trained model from i) (Chung et al., 2023) on FFHQ, and ii) (Dhariwal and Nichol, 2021) on ImageNet. For latent diffusion experiments, we use i) the unconditional LDM-VQ-4 model (Rombach et al., 2022) on FFHQ, and ii) the Stable Diffusion v1.5 (Rombach et al., 2022) model on ImageNet.
@@
We evaluate EquiReg on linear and nonlinear restoration tasks for natural images (see Appendix˜E for task details). We adopt the pre-trained encoder-decoder $\mathcal{E}$-$\mathcal{D}$ as our MPE function. For FFHQ, we use vertical reflection as the symmetry group, which preserves upright facial orientation. For ImageNet, we define a rotation group $G=\{0,\nicefrac{{\pi}}{{2}},\pi,\nicefrac{{3\pi}}{{2}}\}$, and uniformly at random select the group action for each sample. Finally, the loss functions given in Equation˜5 are used to regularize. While our main experiment explore the reflection and rotation groups with small cardinality, EquiReg does not rely on full group coverage. Sampling even a sparse or randomly chosen subset of group actions is sufficient, as long as the function used for regularization exhibits the MPE property across the group (Table˜10).
We characterize EquiReg under change of its period, regularization parameter $\lambda_{t}$, and DDIM steps. EquiReg can be applied efficiently with longer period’ it preserves performance when applied with lower frequency (Table˜1(c)a); and it is robust to the choice of its hyperparameter $\lambda_{t}$, demonstrated on DPS and PSLD (Table˜1(c)b). EquiReg maintains strong performance as the number of DDIM steps is reduced, whereas DPS suffers a significant drop. Importantly, Equi-DPS consistently outperforms DPS, with the performance gap widening at lower step counts (Table˜1(c)c, Figure˜12). An important advantage of EquiReg is that it allows the user to reduce the number of measurement consistency optimization steps, thereby introducing an implicit acceleration in solving the inverse problem. Table˜2 shows that EquiReg regularization enables SITCOM to achieve superior performance with significantly reduced runtime by using fewer measurement consistency steps.
@@
Table˜3a, Table˜4, and Table˜5 highlights the benefits of EquiReg for latent models by consistently improving the performance of ReSample and PSLD across several tasks on FFHQ and ImageNet. We attribute this improvement in part to the reduction of failure cases (Figure˜5d). EquiReg also significantly improves the performance of pixel-based methods (see Equi-DPS vs. DPS, Table˜3d). We observe that EquiReg achieves its largest improvements on perceptual metrics (FID and LPIPS), suggesting it generates more realistic images that lie closer to the data manifold (see Appendix E for supporting qualitative results). EquiReg improves performance under high measurement noise (Figure˜5b). This result aligns with Figure˜5a, which shows the equivariance error is lower on clean images than noisy ones, indicating that EquiReg enforces an effective denoising. Lastly, we note that EquiReg is robust to regularizing hyperparameter $\lambda_{t}$ (Figure˜5c, see Appendix˜C for details). For qualitative performance of EquiReg, see Figure˜15 and Figure˜16.
Table 6: Solving PDEs from sparse observations.
@@
Solving PDEs from sparse observations. EquiReg is evaluated on two important PDE problems: the Helmholtz and Navier-Stokes equations (see Appendix˜F). The objective is to solve both forward and inverse problems in sparse sensor settings. The forward problem involves predicting the solution function or the final state using measurements from only $3\%$ of the coefficient field or the initial state. The inverse problem, conversely, aims to predict the input conditions from observations of $3\%$ of the system’s output. This task is challenging due to the nonlinearity of the equations, the complex structure of Gaussian random fields, and the sparsity of observations.
@@
Recent studies (Huang et al., 2024; Mammadov et al., 2024a; Yao et al., 2025) have demonstrated the superiority of diffusion models over deterministic single-forward methods for solving PDEs. DiffusionPDE (Huang et al., 2024) decomposes the conditional log-likelihood into a learned diffusion prior and a measurement score. FunDPS (Yao et al., 2025) extends the sampling process to a more natural infinite-dimensional spaces, achieving better accuracy and speed via function space models.
@@
We integrate EquiReg into the state-of-the-art FunDPS framework (Mammadov et al., 2024a; Yao et al., 2025), where the equivariance loss is computed using an FNO trained on the corresponding inverse problem. We employ reflection symmetry (i.e., flipping along the $y=x$ axis) and observe no significant performance differences when using alternative transformations such as rotations or alternating flips. Equi-FunDPS improves performance (Table˜6), measured by relative $\ell_{2}$ error, across multiple tasks, particularly in inverse problems where a strong data prior is essential. Importantly, these results highlight that EquiReg is not limited to conventional neural networks and is naturally applicable to neural operators, enabling regularization in both finite-dimensional and function-space diffusion models.
Diversity analysis. To study posterior sampling diversity of EquiReg, we generated $K=10$ posterior samples for $20$ test images across three inverse problems of box inpainting, Gaussian deblurring, $4\times$ super-resolution, and measured diversity using two complementary metrics: Intra-LPIPS for perceptual diversity and Pixel-Std for spatial diversity. Table˜7 demonstrates that Equi-DPS achieves favorable fidelity-diversity trade-offs. We further investigated diversity scaling by varying box inpainting mask size from $128\times 128$ to $192\times 192$ pixels (Figure˜17). Results show that diversity metrics increase linearly with task difficulty, demonstrating that Equi-DPS naturally expands sampling as problems become more ill-posed rather than artificially constraining solutions. This linear relationship indicates healthy, predictable posterior sampling behavior across the difficulty spectrum. Lastly, Figures˜6 and 18 provides qualitative confirmation through visual examples showing four posterior samples per image. Observable variations in facial features, expressions, and eye gaze validate our quantitative measurements, confirming EquiReg can generate genuinely diverse reconstructions rather than collapsing to a single solution (for additional discussion and analysis, see Appendix˜D).
Figure 5: EquiReg effectiveness and robustness across a range of measurement noise levels and regularization parameter (a) Equivariance error computed over a pre-trained decoder on increasingly noisy inputs. (b) EquiReg performance computed over a range of measurement noise levels on the FFHQ dataset. (c) EquiReg reduces failure cases and enhances reconstruction fidelity for super-resolution on FFHQ.
Figure 6: Qualitative diversity examples for super-resolution. We show $K=4$ posterior samples for two test images. Samples differ in facial features (i.e., teeth in the first test image, eye color and eyelashes in the second test image) while maintaining high fidelity to the ground truth, demonstrating EquiReg generates diverse plausible reconstructions rather than collapsing to a single mode.
MPE functions are widespread and effective for EquiReg. We empirically evaluate whether MPE behaviour naturally emerges across commonly used neural networks. We examine i) the emergence of MPE properties in different functions (neural networks) and ii) the effect of using these functions within EquiReg on identical inverse problem settings. For FFHQ and ImageNet, we analyze four function classes: the LDM encoder (Rombach et al., 2022), a CNN autoencoder trained with symmetry augmentations (flip for FFHQ and rotation for ImageNet), pre-trained ResNet-50 (He et al., 2016), and CLIP (Radford et al., 2021). Across all architectures, equivariance error increases systematically as Gaussian noise pushes samples off the data manifold, confirming clear MPE behaviour. The strength of this effect varies: the CNN autoencoder exhibits the strongest MPE signal, while the LDM encoder shows the weakest; ResNet-50 and CLIP lie between these extremes (see Figures˜14 and 19). Notably, the strongest MPE behaviour of the CNN autoencoder is in line with our systematic guidelines for constructing MPE functions; This is precisely the regime where the training distribution of the function matches the distribution of the inverse problem (i.e., training on ImageNet train and evaluating on ImageNet test).
We then integrate each function into EquiReg under identical inverse problem settings (FFHQ/ImageNet, DPS/SITCOM, super-resolution/motion deblurring). In all cases, EquiReg improves reconstruction quality over the corresponding baseline without regularization (Tables˜8 and 13). These results demonstrate that i) MPE behaviour naturally arises in widely used pre-trained networks, and ii) EquiReg is robust to the choice of MPE function, even when the MPE signal is relatively weak. Notably, our main experiments use the weakest MPE (the LDM encoder), suggesting further gains are possible with stronger MPE constructions.
Text-to-image guidance. Given a “source” image, DreamSampler (Kim et al., 2024) transforms it according to a text prompt. When applying EquiReg to DreamSampler, perceptual quality and reduced artifacts are improved in the generated images. Figure˜2 shows a source “cat” transformed into a “corgi”: Equi-DreamSampler produces more realistic results and resolves anatomical inconsistencies (e.g., correcting a three-front-legged corgi to two front legs). We also observe an implicit acceleration effect when EquiReg is imposed (Figure˜7). Equi-DreamSampler with $50$ DDIM steps produces images comparable to DreamSampler with substantially more steps. Increasing the regularization strength $\lambda_{t}$ at fixed $50$ DDIM steps results in effects similar to increasing DDIM steps in DreamSampler (from $50$ to $75$ to $100$), suggesting that EquiReg guides sampling trajectories closer to the data manifold. These experiments illustrate that EquiReg is readily applicable beyond inverse problems, including text-to-image guidance. While not our primary focus, they demonstrate the broader utility and flexibility of EquiReg as a plug-and-play regularization framework.
Table 7: Fidelity and diversity comparison across inverse problems. EquiReg improves fidelity while largely preserving or enhancing sampling diversity (2020 test images with $K=10$ samples per image.)
Task Method Fidelity Metrics Diversity Metrics
[…]
Box inpainting DPS 0.140 70.89 0.112 9.286
Equi-DPS (ours) 0.112 59.70 0.118 10.59
Gaussian deblur DPS 0.150 76.71 0.114 6.565
Equi-DPS (ours) 0.120 63.02 0.092 5.669
Super-resolution ($\times 4$) DPS 0.683 99.11 0.134 7.956
Equi-DPS (ours) 0.703 87.52 0.187 23.52
Table 8: MPE behaviour emerges across diverse network architectures and consistently improves diffusion-based inverse problems with EquiReg. SITCOM motion deblurring on FFHQ 256 with $\lambda=0.05$. Results are reported as mean (standard deviation).
MPE function PSNR SSIM LPIPS
None 27.670 (1.343) 0.790 (0.031) 0.221 (0.040)
LDM Encoder (FFHQ) 28.357 (1.379) 0.806 (0.031) 0.200 (0.036)
CNN Autoencoder (FFHQ) 28.852 (1.376) 0.819 (0.044) 0.193 (0.033)
Pretrained ResNet50 28.682 (1.388) 0.811 (0.036) 0.198 (0.036)
@@ 5 Conclusion
We introduce Equivariance Regularized (EquiReg) diffusion for inverse problems. EquiReg regularizes sampling trajectories to stay closer to the data manifold, leveraging manifold-preferential equivariance (MPE): functions with low equivariance error on-manifold and high error off-manifold. Such functions arise naturally in trained networks and can serve as plug-and-play regularizers without modifying the diffusion denoiser. EquiReg is agnostic across pixel- and latent-space diffusion models and remains robust under reduced sampling, effectively accelerating convergence. Across diverse inverse problems, it consistently improves perceptual and reconstruction metrics while reducing failure cases, highlighting its generality and efficiency.
EquiReg operates as a plug-and-play regularization framework and therefore builds upon the quality of the underlying diffusion solver. It does not modify the diffusion architecture itself, but instead improves sampling by guiding trajectories toward more consistent, on-manifold solutions. As a regularization mechanism, its impact is most pronounced in challenging regimes where likelihood guidance alone may degrade or become unstable. Applying EquiReg requires selecting appropriate symmetry groups and constructing suitable MPE functions for the task at hand. While we provide systematic guidelines for imaging and PDE settings, extending these constructions to new domains remains an important direction for future work; we have provided reference to prior work on how symmetries for various application can be learned and set. Finally, although we formalize distribution-dependent equivariant and MPE functions, a deeper theoretical understanding of when MPE behaviour emerges in trained networks and how it may interact with joint training of diffusion models, remains an interesting avenue for further study.
Broader Impact Statement. On the positive side, high‑fidelity image restoration can improve downstream tasks in medical imaging, remote‑sensing and environmental monitoring (e.g., denoising satellite observations to track pollution or deforestation). Likewise, accelerated PDE‑solving via learned diffusion priors may enable faster, more accurate simulations for climate modeling, fluid‑dynamics research, and engineering design. On the other hand, robust reconstruction methods could be misappropriated for privacy‑invasive surveillance or to create deceptive imagery. We emphasize that our method does not amplify these existing risks.
Appendices for “EquiReg: Equivariance Regularized Diffusion for Inverse Problems”
@@
We provide our source code when EquiReg. We will provide a publicly available source code upon acceptance. This supplementary materials contain the following:
@@
  • Appendix˜A includes additional experiments on text-to-image guidance. We regularize DreamSampler (Kim et al., 2024) with EquiReg for an improved performance (see Figures˜7, 8, 9, 10 and 11).
  • Appendix˜B includes additional experiments on robustness including robustness to $\lambda_{t}$, reduced number of DDIM steps, and reduced number of measurent consistency steps.
@@
  • Appendix˜C includes qualitative analysis on the performance of methods with and without EquiReg. Results show a reduction of artifacts and an improved perceptual quality of the solution. This section also includes the equivariance error of a pre-trained encoder used in EquiReg (Figure˜14(a)) and a histogram of Equi’s improvement on DPS (Figure˜13).
  • Appendix˜D includes diversity experiments. Results show that EquiReg achieves favorable fidelity-diversity tradeoffs (Table˜7, Figure˜17, and Figure˜18).
@@
  • Appendix˜E demonstrates EquiReg experimental setup and implementation for PSLD, ReSample, and DPS (Algorithms˜2, 3, 4, 5 and 6). It also contains information about the EquiReg hyperparameters for image restoration tasks.
@@
  • Appendix˜F contains information on the PDE reconstruction experiment. It discusses the equations along with implementation details and hyperparameters.
  • Appendix˜G provides theoretical proofs of Propositions˜G.1 and G.2.
@@
  • Appendix˜H contains additional background information on solving inverse problems, vanishing-error autoencoders, and equivariance.
  • Appendix˜I provides additional figures on MPE functions.
  • Appendix˜J discloses computing resources used to conduct the experiments.
  • Appendix˜K credits code assets used for our experiments.
  • Appendix˜L concludes the appendix with a “responsible release” statement.
The authors acknowledge the usage of LLMs on proofreading and improving the coherency of the manuscript. The authors have not used LLMs for content generation.